Two cars start from the same point and travel in opposite directions with speeds of 45 and 60 miles per hour. In how many hours will they be 490 miles apart?
Found 2 solutions by rfer, Jk22:
Answer by rfer(16322)

moving apart at 105 mph
490/105=4.667=4 hrs 20 min

Answer by Jk22(389)

45 and 60 equals 105mph is the speed the cars go away from each other.

Then 105*t=490, t=490/105=4hours2/3=4hours and 40 minutes.